ok here is a question that is troubling me. the directions say "determine the value of "r" so that a line through the points with the given coordinates has the given slope. draw a sketch of each situation" now about this sketch thing i dont think i need help on that you dont got to worry about that. but here is the problem. (5,r), (2,3); slope=2. thank you for taking the time to help me.
Answer by longjonsilver(2297)

given 2 points, the gradient is (change in y)/(change in x) and this equals 2.

so, (r-3)/(5-2) = 2
(r-3)/3 = 2
r-3 = 6
r = 9

jon.